Warren Company plans to depreciate a new building using the double declining-balance depreciation method. The building cost is $700,000. The estimated residual value of the building is $40,000 and it has an expected useful life of 10 years.

What is the building’s book value at the end of the first year?
Answer #1

Double declining rate = 200/Useful life = 200/10 = 20%

Depreciation expense for first year = 700,000*20% = 140,000

Book value = Cost - Depreciation expense for the first year

= 700,000 - 140,000

= 560,000